What is the equation in slope intercept form for a line with a slope of 7/3 and passes through the points (7,0) & (10,7)?
What are their x and y-intercepts? What are their (x,y) coordinates?

Answer by stanbon(75887)

What is the equation in slope intercept form for a line with a slope of 7/3 and passes through the points (7,0) & (10,7)?
---
slope = 7/3
Form: y = mx + b
0 = (7/3)*7 + b
0 = (49/3) + b
b = (-49/3)
-----
Equation:
y = (7/3)x-(49/3)
==========================
What are their x and y-intercepts? What are their (x,y) coordinates?
x-int: Let y = 0, then x = 7 as was shown in the point (7,0)
---
y-int: let x = 0, then y = -49/3 giving you (0, -49/3)
======================================
Cheers,
Stan H.